OpenStudy (anonymous):

What is the product in lowest terms? 8/15 * -3/10 A. -4/25 B. -3/35 C. - 2/25 D. 1/5 Let me solve the problem myself and you'll tell me if it's correct

OpenStudy (mathstudent55):

Read the similar problem I just solved a few more times until you understand it. Then follow the same procedure and try to solve this one. Post what you get. If you still have questions, ask.

OpenStudy (anonymous):

\[\frac{ 8 }{ 15 } \times \frac{ -3 }{ 10 } = \frac{ -24 }{ 150 }\]

OpenStudy (mathstudent55):

very good now we need to reduce

OpenStudy (mathstudent55):

what are 24 and 150 both divisible by?

OpenStudy (anonymous):

wait a sec.....I'm reducing

OpenStudy (anonymous):

THE FINAL ANSWER IS \[\frac{ -4 }{ 25 }\]

OpenStudy (anonymous):

SO IT'S A

OpenStudy (mathstudent55):

yes

OpenStudy (anonymous):

THANKS

OpenStudy (mathstudent55):

Write is as \(-\dfrac{4}{25} \)

OpenStudy (texaschic101):

 yep....try using the shortcut